The sequence 2, 9, 16, 23, 30,.... is an arithmetic progression. Find the 33rd term in the progression. A/
Answer #1

2,9,16,23,3 0, nith term of airthmetic progression Tn = a + (n-1)d where a is furnt term d is common different a. 2 d=9-2=7 3
Answer #2
-12,-5,2,9,15
source: Algebra sequence
answered by: Jay
Answer #3
-12,-5,2,9,15
source: Algebra sequence
answered by: Jay